# EfficientQAT: Efficient Quantization-Aware Training for Large Language Models

### v1

@@ Abstract
Large language models (LLMs) are integral to modern natural language processing and artificial intelligence. However, they face challenges in managing their significant memory requirements. Although quantization-aware training (QAT) offers a solution by reducing memory consumption through low-bit representations with minimal accuracy loss, it demands substantial training resources to optimize model weights and quantization parameters. To address this, we propose Efficient Quantization-Aware Training (EfficientQAT), a novel quantization technique for compressing LLMs. EfficientQAT involves two consecutive phases: Block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). Block-AP sequentially conducts quantization-aware training for all parameters in each transformer block with block-wise reconstruction, maintaining efficiency by avoiding training the entire LLM. Initialized with quantized model, E2E-QP then trains only quantization parameters (step sizes) end-to-end, enhancing efficiency with a fixed quantized backbone and reduced trainable parameter count. Extensive experiments demonstrate that EfficientQAT outperforms previous quantization methods across a range of models, including base LLMs, instruction-tuned LLMs, and multimodal LLMs, with scales from 7B to 70B parameters at various quantization bits. For instance, EfficientQAT obtains a 2-bit Llama-2-70B model on a single A100-80GB GPU in 41 hours, with less than 3% accuracy degradation compared to the full precision (69.48 vs. 72.41). Notably, this INT2 quantized 70B model obtains a 1.67 accuracy gain over the Llama-2-13B model (69.48 vs. 67.81) while requiring less memory (19.2GB vs. 24.2GB). Code is available at https://github.com/OpenGVLab/EfficientQAT.
@@ 1 Introduction
Recent advancements in large language models (LLMs) llama2 ; gpt4 ; vicuna ; lvlm-ehub ; mmtbench have demonstrated impressive capabilities in diverse language tasks such as reasoning arc ; boolq ; hellaswag , cognitive processing mme ; lvlm-ehub , and agent-based applications tool_learning ; toolllm . However, these models are characterized by their extensive parameters, which pose significant challenges for memory footprint and bandwidth squeezellm ; besa .
Quantization-aware training (QAT), one of the most effective quantization techniques, works by minimizing quantization errors through training with quantization constraints. Although QAT can compress LLMs effectively without significant performance loss, it requires training the whole LLM on a large corpus, resulting in enormous training costs. For instance, the QAT method BitNet b1.58 bitnet158 can achieve nearly lossless ternary quantization but requires retraining LLMs from scratch using the full pre-trained dataset, which is impractical for extremely large models***BitNet b1.58 only verifies the performance on 3B models with 100B training tokens..
In pursuit of efficient quantization for large language models (LLMs), techniques such as post-training quantization (PTQ) and quantized parameter-efficient fine-tuning (Q-PEFT) have been developed. PTQ awq ; gptq ; omniquant ; quip ; aqlm minimizes memory footprint during inference by converting pre-trained LLM weights from 16-bit to formats like 2-bit without retraining. While PTQ is suitable for quick deployment with minimal resources, it may compromise accuracy, especially in low-bit regimes, as shown in Figure 1(a). Q-PEFT qlora ; qa-lora ; loftq ; lq-lora ; peqa , on the other hand, uses parameter-efficient fine-tuning methods like LoRA lora to train quantized models, significantly reducing memory requirements during training. For instance, QLoRA qlora converts the model from FP16 to NF4 (a 4-bit float format) and then applies a LoRA module, allowing fine-tuning on consumer-grade GPUs without sacrificing performance. However, most Q-PEFT techniques qlora ; lq-lora ; loftq revert to FP16 for integrating additional LoRA modules post fine-tuning. This necessitates another round of PTQ for deployment on memory-limited platforms, which can degrade performance. Although solutions like PEQA peqa and QA-LoRA qa-lora attempt to address this issue, their performance remains suboptimal in low-bit scenarios (2-bit and 3-bit) due to poor quantized model initialization (Figure 1(b)).
@@
Figure 1: (a) EfficientQAT significantly surpasses existing uniform quantization methods, and is either superior to or comparable with vector quantization techniques. (b) EfficientQAT markedly outperforms existing Q-PEFT methods.
(a) inference speed comparisons
(b) training speed comparisons
Figure 2: (a): This panel shows a speed comparison of vector-matrix multiplications for the gate_proj component of Llama-2-70B. Uniform Quantization proves more hardware-efficient than vector quantization. (b): EfficientQAT can complet the QAT of 70B models on a single A100 GPU with less time.
To tackle the above challenges, we introduce a novel quantization technique named EfficientQAT, to compress LLMs effectively and efficiently. EfficientQAT surpasses existing PTQ omniquant ; aqlm , Q-PEFT qa-lora ; qlora and QAT llmqat methods in performance while maintaining memory efficiency during training and inference, as illustrated in Figure 2(a) & 2(b). To improve quantization efficiency and accuracy, EfficientQAT involves two consecutive phases: Block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). Block-AP sequentially conducts quantization-aware training for all parameters including the original full-precision weights and quantization parameters (step sizes and zero points) in each transformer block with block-wise reconstruction. While Block-AP trains all parameters, it preserves memory efficiency by avoiding training the entire LLM. Moreover, potential overfitting issues in Block-AP can be mitigated by simply increasing the training samples from 128 to 4096 (see Figure 5). Notably, our fully trainable paradigm of Block-AP outperforms other trainable variants (see Table 7), such as training rounding adaround ; weight_round , clipping thresholds omniquant , and step sizes lsq ; cbq .
Building on the robust initialization provided by Block-AP, E2E-QP fixes the quantized weights and trains only the quantization parameters (step sizes) end-to-end, further enhancing the quantization performance. Moreover, E2E-QP enables instruction tuning of quantized models for specific tasks in a quantization-aware and memory-efficient manner, as trainable quantization parameters take only a small fraction of the total network (e.g., 1.5% with a quantization group size of 64). It works like previous quantized parameter-efficient fine-tuning (Q-PEFT)techniques such as QA-LoRA qa-lora . Therefore, EfficientQAT is general enough to compress various models including base LLMs (Llama llama2 ; llama , instruction-tuned LLMs (Vicuna vicuna ), and multimodal LLMs (LLaVA llava1.5 ).
@@
Thanks to the integration of the Block-AP and E2E-QP, EfficientQAT characterizes itself as a fast-converging, memory-efficient, and high-performing quantization technique. For instance, EfficientQAT can obtain a 2-bit Llama-2-70B model on a single A100-80GB GPU in just 41 hours, with less than 3% zero-shot accuracy degradation compared to its full-precision counterpart (69.48 vs. 72.41). We also evaluate EfficientQAT across scenarios involving model compression and instruction-tuning. In model compression, as illustrated in Figure 1(a), EfficientQAT significantly outperforms other uniform quantization methods by approximately 5 points on accuracy in the challenging 2-bit quantization setting. It also matches the performance of vector quantization methods aqlm ; quip-sharp in this scenario. Furthermore, Figure 2(a) demonstrates that the uniform quantization implemented by EfficientQAT is more hardware-efficient than vector quantization, which introduces a considerable overhead. In terms of instruction tuning, as shown in Figure 1(b), EfficientQAT consistently outperforms existing Q-PEFT methods, including QLoRA qlora , QA-LoRA qa-lora , and PEQA peqa .
@@ 2 Related Works
Post-Training Quantization of LLMs. PTQ is a pivotal technique for accelerating and deploying LLMs. Quantization approaches generally fall into two categories: weight-only quantization gptq ; spqr ; owq ; squeezellm ; norm_tweaking ; teq and weight-activation quantization smoothquant ; qllm ; os ; os+ ; rptq ; atom ; quik ; odyssey ; quarot . Weight-only quantization focuses on compressing weights into low-bit formats, reducing memory demands and enhancing the efficiency of memory-bounded computations in LLMs qserve ; llm-unveil . Conversely, weight-activation quantization compresses both weights and activations, thus further decreasing the overhead associated with matrix multiplications qserve . Recent advancements in weight-only quantization include the introduction of vector quantization methods by QUIP#quip-sharp and AQLMaqlm . These methods have shown promising performance but also introduce significant overhead llm-qbench . Our research continues to explore uniform quantization, which is preferred for its compatibility with hardware implementations.
Quantization-Aware Training of LLMs. QAT can enhance the performance of quantized models beyond what PTQ offers. However, QAT has been less explored in LLMs due to the significant training costs involved. Studies such as LLM-QAT llmqat and BitDistiller bitdistiller investigate the application of knowledge distillation within QAT contexts. Techniques like BitNet b1.58 bitnet158 and OneBit onebit employ QAT to achieve extreme binary or ternary quantization levels. Although BitNet b1.58 demonstrates near-lossless performance on models up to 3 billion parameters and 100 billion training tokens with ternary quantization, its applicability to larger models or datasets remains uncertain due to prohibitive training expenses.
Quantized Parameter-Efficient Fine-Tuning of LLMs. Techniques like QLoRA qlora , INT2.1 int2 , LQ-LoRA lq-lora , and LoftQ loftq quantize model parameters to low-bit representations followed by the addition of LoRA lora modules for fine-tuning. However, these methods require merging the LoRA modules into quantized weights, resulting in the model reverting to the FP16 format. Addressing this issue, QA-LoRA qa-lora redesign the LoRA module to merge seamlessly into the zero points. To our knowledge, the work most closely to our approach is PEQA peqa , which employs a simple round-to-nearest (RTN) method for low-bit quantization and fine-tunes the step sizes for downstream task adaptation. Despite this, the suboptimal PTQ initialization technique utilized by PEQA leads to performance that falls short when compared to QLoRA qlora and QA-LoRA qa-lora .
@@ 3.1 Method Overview
In this section, we introduce EfficientQAT, a novel quantization-aware training framework for LLMs that enhances memory efficiency. As illustrated in Figure 3, traditional QAT approaches train the full-precision weights $\mathbf{W}$ and quantization parameters $s$ (step sizes) and $z$ (zero points) simultaneously in an end-to-end manner, which significantly increases the memory requirements due to the large number of parameters involved. To address this issue, EfficientQAT adopts a two-stage strategy: block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). In the Block-AP phase, model parameters and quantization parameters are trained block-by-block using reconstruction loss, which not only allows for precise calibration with full training but also reduces memory consumption brecq ; omniquant by block-wise training. Following this, the E2E-QP phase fixes the quantized weights and trains the step sizes exclusively on target datasets, thus achieving quantization-aware training in a memory-efficient way. Details on Block-AP and E2E-QP are further described in Sections 3.2 and 3.3, respectively.
@@ 3.2 Block-Wise Training of All Parameters
Blcok-wise Quantization-aware Training. Traditional QAT methods bitnet158 ; lsq ; llmqat train the entire network using Eq.(1) and Eq.(2) in an end-to-end fashion, which typically requires substantial computational resources and extensive data to prevent overfitting. Here we aim to enhance the training efficiency of QAT. Previous studies, such as BRECQ brecq , have demonstrated that block-wise training achieves faster convergence and requires less training time, data, and memory than end-to-end training given a pre-trained model. Following the methodologies in BRECQ brecq and OmniQuant omniquant , Block-AP sequentially conducts quantization-aware training within one transformer block before moving on to the next under a block-wise reconstruction framework.
Full Training of Model Weights and Quantization Parameters. Unlike previous methods which optimize several quantization parameters such as rounding parameters adaround ; weight_round ; flexround , clipping parameters omniquant , and step sizes lsq ; cbq , Block-AP behaves like QAT, training all inherent parameters from Eq.(1) and Eq.(2), including scaling factor $s$, zero point $z$, and model weights $\mathbf{W}$.
In our Block-AP approach, a straightforward full-training regimen outperforms existing partial-training variants adaround ; brecq ; cbq with intricate designs. Traditional training methods involving rounding parameters adaround ; brecq ; cbq serve as regularization techniques, constraining the update range of integral weights to $(-1,+1)$ to mitigate overfitting. However, this approach limits the solution space, potentially hindering the final performance of quantized models. By simply scaling the training samples from $128$ to $4096$, Block-AP effectively addresses the overfitting issue (see Figure 5 for details) while improving performance compared to training with rounding parameters. Our empirical findings demonstrate the superiority of full training within our Block-AP over existing partial-training variants adaround ; brecq ; cbq , as shown in Table 7.
@@ 3.3 End-to-End Training of Quantization Parameters
End-to-End Training of step sizes. Unlike traditional Quantization-Aware Training (QAT) methods llmqat ; bitnet158 that train full-precision weights, E2E-QP begins with $\mathbf{W}_{q}$ initialized via Block-AP and focuses solely on the training of quantization parameters ($s$ and $z$). Our findings indicate that training $s$, $z$, or both yields similar performance (see Table 7 for details). However, since training $z$ involves converting it from a low-bits format to full-precision, we typically train only $s$ by default unless specified otherwise to avoid additional memory overhead.
@@
Overall, the memory usage for training in E2E-QP is drastically reduced due to the reduced trainable parameter count. Detailed memory footprints for various model sizes and bits under E2E-QP are listed in Table 8. For instance, the Llama-2-70B model can complete 2-bit QAT through E2E-QP using only 34.2GB of memory. Equipped with E2E-QP, EfficientQAT is adaptable to different scenarios by simply changing the training datasets, which includes applications such as continual pre-training and instruction-tuning alpaca .
Table 1: Llama 2 & 3 average zero-shot accuracy on 5 common-sense reasoning tasks ($\uparrow$). Detailed number of each task can be found at Table 13 (4-bit), Table 14 (3-bit), and Table 15 (2-bit). “Group” indicates group size for uniform quantization and codebook scheme for vector quantization.

### v2

@@ Abstract
Large language models (LLMs) are crucial in modern natural language processing and artificial intelligence. However, they face challenges in managing their significant memory requirements. Although quantization-aware training (QAT) offers a solution by reducing memory consumption through low-bit representations with minimal accuracy loss, it is impractical due to substantial training resources. To address this, we propose Efficient Quantization-Aware Training (EfficientQAT), a more feasible QAT algorithm. EfficientQAT involves two consecutive phases: Block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). To the best of our knowledge, Block-AP is the first method to enable direct training of all parameters in a block-wise manner, reducing accuracy loss in low-bit scenarios by enhancing the solution space during optimization. E2E-QP then trains only the quantization parameters (step sizes) end-to-end, further improving the performance of quantized models by considering interactions among all sub-modules. Extensive experiments demonstrate that EfficientQAT outperforms previous quantization methods across a range of models, including base LLMs, instruction-tuned LLMs, and multimodal LLMs, with scales from 7B to 70B parameters at various quantization bits. For instance, EfficientQAT obtains a 2-bit Llama-2-70B model on a single A100-80GB GPU in 41 hours, with less than 3 points accuracy degradation compared to the full precision (69.48 vs. 72.41). Code is available at https://github.com/OpenGVLab/EfficientQAT.
@@ 1 Introduction
Recent advancements in large language models (LLMs) (Touvron et al., 2023; Bubeck et al., 2023; Chiang et al., 2023; Xu et al., 2023a; Ying et al., 2024) have demonstrated impressive capabilities in diverse language tasks such as reasoning (Clark et al., 2018; 2019; Zellers et al., 2019), cognitive processing (Fu et al., 2023; Xu et al., 2023a), and agent-based applications (Qin et al., 2023a; b). However, these models are characterized by their extensive parameters, which pose significant challenges for memory footprint and bandwidth (Kim et al., 2023b; Xu et al., 2024a).
Quantization-aware training (QAT) is a highly effective quantization technique that minimizes quantization errors by incorporating quantization constraints during training. For example, BitNet b1.58 (Ma et al., 2024) can achieve nearly lossless ternary quantization. The precision of QAT is due to two main factors: 1) Fully trainable parameters allow for enough optimized space for gradient descent optimization; 2) End-to-end training accounts for interactions among all sub-modules in the models. Despite its performance benefits, QAT demands significant training resources, such as time and GPUs, as well as extensive training data. For instance, BitNet b1.58 requires retraining LLMs from scratch using the entire pre-trained dataset. Therefore, this approach is impractical for extremely large models and has only been verified on 3B models with 100B training tokens.
In optimizing quantization for LLMs, current methods emphasize either fine-grained reconstruction or reducing trainable parameters. While these approaches improve efficiency, they significantly degrade accuracy in low-bit scenarios. Mainstream post-training quantization (PTQ) methods (Lin et al., 2023; Frantar et al., 2022; Shao et al., 2023) focus on block-wise reconstruction (Li et al., 2021). They also restrict the optimization space to alleviate overfitting risk by only training rounding parameters (Nagel et al., 2020; Cheng et al., 2023), clipping thresholds (Shao et al., 2023), or step sizes (Esser et al., 2019; Ding et al., 2023). However, these methods not only limit optimizable parameters but also overlook cross-block interactions, leading to notable accuracy degeneration in low-bit scenarios, as shown in Figure 1(a). Conversely, quantized parameter-efficient fine-tuning (Q-PEFT) methods (Dettmers et al., 2023a; Kim et al., 2023a) reduce training costs by freezing quantized parameters and only training a few continuous floats. For example, PEQA (Kim et al., 2023a) and QA-LoRA (Xu et al., 2023b) focus on training continuous quantization parameters. Despite this, their performance remains poor, as depicted in Figure 1(b), because the severe performance loss in low-bit scenarios (2-bit and 3-bit) cannot be fully recovered with limited trainable parameters.
@@
Figure 1: (a) EfficientQAT significantly surpasses existing uniform quantization methods, and is either superior to or comparable with vector quantization techniques. (b) EfficientQAT markedly outperforms existing Q-PEFT methods. (c) EfficientQAT can complete the QAT of 70B models on a single NVIDIA A100-80GB GPU with less time.
To address these challenges, we introduce a novel quantization-aware training framework called EfficientQAT. This framework combines the advantages of fully trainable parameters and end-to-end training, similar to native QAT (Ma et al., 2024), while maintaining the training efficiency of PTQ (Cheng et al., 2023; Shao et al., 2023) and Q-PEFT (Xu et al., 2023b). EfficientQAT introduces block-wise training of all parameters (Block-AP) to enhance the optimizable space and mitigate quantization accuracy loss. Block-AP sequentially trains all parameters, including original full-precision weights and quantization parameters (step sizes and zero points), within each transformer block. Several works have been developed based on block-wise reconstruction. However, previous approaches focus on designing additional trainable parameters, such as clipping thresholds for OmniQuant (Shao et al., 2023), weight rounding for AutoRound (Cheng et al., 2023) and BRECQ (Li et al., 2021), or LoRA (Hu et al., 2021) parameters for CBQ (Ding et al., 2023). Our Block-AP is the first to directly train all parameters during block-wise reconstruction, achieving superior performance compared to previous methods (see Table 7). Block-AP successfully demonstrates that complex trainable parameter design is unnecessary for effective block-wise reconstruction in LLMs quantization. Furthermore, we introduce end-to-end training of quantization parameters (E2E-QP) to account for inter-block interactions. E2E-QP keeps the quantized weights fixed and trains only the quantization parameters (step sizes) end-to-end.
@@
Thanks to the integration of the proposed Block-AP and E2E-QP, EfficientQAT characterizes itself as a fast-converging, memory-efficient, and high-performing quantization technique. For instance, EfficientQAT can obtain a 2-bit Llama-2-70B model on a single A100-80GB GPU in just 41 hours, with less than 3 points accuracy degradation on 5 zero-shot common-sense tasks compared to its full-precision counterpart (69.48 vs. 72.41). We also evaluate EfficientQAT across scenarios involving model compression and instruction-tuning. In model compression, as illustrated in Figure 1(a), EfficientQAT significantly outperforms existing uniform quantization methods by approximately 5 points on accuracy in the challenging 2-bit quantization setting. It also matches the performance of vector quantization methods (Egiazarian et al., 2024; Tseng et al., 2024) in this scenario. In terms of instruction tuning, as shown in Figure 1(b), EfficientQAT consistently outperforms existing Q-PEFT methods, including QLoRA (Dettmers et al., 2023a), QA-LoRA (Xu et al., 2023b), and PEQA (Kim et al., 2023a). For instance, EfficientQAT surpasses PEQA (Kim et al., 2023a) with 4.5 points MMLU accuracy when fine-tuning with Alpaca dataset.
@@ 2 Related Works
Post-Training Quantization of LLMs. PTQ is a pivotal technique for accelerating and deploying LLMs. Quantization approaches generally fall into two categories: weight-only quantization (Frantar et al., 2022; Dettmers et al., 2023b; Lee et al., 2023a; Kim et al., 2023b) and weight-activation quantization (Xiao et al., 2023; Liu et al., 2023c; Wei et al., 2022; 2023; Yuan et al., 2023; Zhao et al., 2023; Ashkboos et al., 2023; Li et al., 2023a; Ashkboos et al., 2024). Weight-only quantization focuses on compressing weights into low-bit formats, reducing memory demands and enhancing the efficiency of memory-bounded computations in LLMs (Lin et al., 2024; Yuan et al., 2024). Conversely, weight-activation quantization compresses both weights and activations, thus further decreasing the overhead associated with matrix multiplications (Lin et al., 2024). Recent advancements in weight-only quantization include the introduction of vector quantization methods by QUIP#Tseng et al. (2024) and AQLMEgiazarian et al. (2024). These methods have shown promising performance but also introduce significant overhead (Gong et al., 2024). Our research continues to explore uniform quantization, which is preferred for its compatibility with hardware implementations.
Quantization-Aware Training of LLMs. QAT can enhance the performance of quantized models beyond what PTQ offers. However, QAT has been less explored in LLMs due to the significant training costs involved. Studies such as LLM-QAT (Liu et al., 2023e) and BitDistiller (Du et al., 2024) investigate the application of knowledge distillation within QAT contexts. Techniques like BitNet b1.58 (Ma et al., 2024) and OneBit (Xu et al., 2024b) employ QAT to achieve extreme binary or ternary quantization levels. Although BitNet b1.58 demonstrates near-lossless performance on models up to 3 billion parameters and 100 billion training tokens with ternary quantization, its applicability to larger models or datasets remains uncertain due to prohibitive training expenses.
Quantized Parameter-Efficient Fine-Tuning of LLMs. Techniques like QLoRA (Dettmers et al., 2023a), INT2.1 (Chai et al., 2023), LQ-LoRA (Guo et al., 2023), and LoftQ (Li et al., 2023b) quantize model parameters to low-bit representations followed by the addition of LoRA (Hu et al., 2021) modules for fine-tuning. However, these methods require merging the LoRA modules into quantized weights, resulting in the model reverting to the FP16 format. Addressing this issue, QA-LoRA (Xu et al., 2023b) redesigns the LoRA module to merge seamlessly into the zero points. The approach most similar to ours is PEQA (Kim et al., 2023a), which uses a round-to-nearest (RTN) method for low-bit quantization and fine-tunes step sizes for task adaptation. However, PEQA experiences significant performance degradation due to limited trainable parameters, which hinders recovery from quantization information loss.
@@ 3.1 Method Overview
In this section, we introduce EfficientQAT, a novel quantization-aware training framework for LLMs that enhances memory efficiency. As illustrated in Figure 2, traditional QAT approaches train the full-precision weights $\mathbf{W}$ and quantization parameters $s$ (step sizes) and $z$ (zero points) simultaneously in an end-to-end manner, which significantly increases the memory requirements due to the large number of parameters involved. To address this issue, EfficientQAT adopts a two-stage strategy: block-wise training of all parameters (Block-AP) and end-to-end training of quantization parameters (E2E-QP). In the Block-AP phase, model parameters and quantization parameters are trained block-by-block using reconstruction loss, which not only allows for precise calibration with full training but also reduces memory consumption (Li et al., 2021; Shao et al., 2023) by block-wise training. Following this, the E2E-QP phase fixes the quantized weights and trains the step sizes exclusively on target datasets, thus achieving inter-block interaction in a memory-efficient way. Details on Block-AP and E2E-QP are further described in Sections 3.2 and 3.3, respectively.
@@ 3.2 Block-Wise Training of All Parameters
Blcok-wise Quantization-aware Training. Traditional QAT methods (Ma et al., 2024; Esser et al., 2019; Liu et al., 2023e) train the entire network using Eq.(1) and Eq.(2) in an end-to-end fashion, which typically requires substantial computational resources and extensive data to prevent overfitting. Here we aim to enhance the training efficiency of QAT. Previous studies, such as BRECQ (Li et al., 2021), have demonstrated that block-wise training achieves faster convergence and requires less training time, data, and memory than end-to-end training given a pre-trained model. Following the methodologies in BRECQ (Li et al., 2021) and OmniQuant (Shao et al., 2023), Block-AP sequentially conducts quantization-aware training within one transformer block before moving on to the next under a block-wise reconstruction framework.
Full Training of Model Weights and Quantization Parameters. Unlike previous methods which optimize several quantization parameters such as rounding parameters (Nagel et al., 2020; Cheng et al., 2023; Lee et al., 2023b), clipping parameters (Shao et al., 2023), and step sizes (Esser et al., 2019; Ding et al., 2023), Block-AP behaves like QAT, training all inherent parameters from Eq.(1) and Eq.(2), including scaling factor $s$, zero point $z$, and model weights $\mathbf{W}$.
In our Block-AP approach, a straightforward full-training regimen outperforms existing partial-training variants (Nagel et al., 2020; Li et al., 2021; Ding et al., 2023) with intricate designs. Traditional training methods involving rounding parameters (Nagel et al., 2020; Li et al., 2021; Ding et al., 2023) serve as regularization techniques, constraining the update range of integral weights to $(-1,+1)$ to mitigate overfitting. However, this approach limits the solution space, potentially hindering the final performance of quantized models. Our empirical findings demonstrate the superiority of full training within our Block-AP over existing partial-training variants (Nagel et al., 2020; Li et al., 2021; Ding et al., 2023), as shown in Table 7.
@@ 3.3 End-to-End Training of Quantization Parameters
End-to-End Training of step sizes. Unlike traditional Quantization-Aware Training (QAT) methods (Liu et al., 2023e; Ma et al., 2024) that train full-precision weights, E2E-QP begins with $\mathbf{W}_{q}$ initialized via Block-AP and focuses solely on the training of quantization parameters ($s$ and $z$). Our findings indicate that training $s$, $z$, or both yields similar performance (see Table 7 for details). However, since training $z$ involves converting it from a low-bits format to full-precision, we typically train only $s$ by default unless specified otherwise to avoid additional memory overhead.
@@
Overall, the memory usage for training in E2E-QP is drastically reduced due to the reduced trainable parameter count. Detailed memory footprints for various model sizes and bits under E2E-QP are listed in Table 8. For instance, the Llama-2-70B model can complete 2-bit QAT through E2E-QP using only 34.2GB of memory. Equipped with E2E-QP, EfficientQAT is adaptable to different scenarios by simply changing the training datasets, which includes applications such as continual pre-training and instruction-tuning (Taori et al., 2023).
Table 1: Llama 2 & 3 average zero-shot accuracy on 5 common-sense reasoning tasks ($\uparrow$). Detailed number of each task can be found at Table 15 (4-bit), Table 16 (3-bit), and Table 17 (2-bit). “Group” indicates group size for uniform quantization and codebook scheme for vector quantization. “bold” indicates best results for uniform quantization.